OpenStudy (anonymous):

Factor the following polynomial: x2 + 14x – 32 A. (x – 2)(x – 7) B. (x + 7)(x + 2) C. (x + 16)(x – 2) D. (x + 2)(x –16)

OpenStudy (diyadiya):

Do you know to factorise?

OpenStudy (anonymous):

NO

OpenStudy (diyadiya):

We have to split the middle term for that you need two numbers whose product is -32 and sum is 14 what can those two numbers be?

OpenStudy (anonymous):

D ?

OpenStudy (anonymous):

@Diyadiya

OpenStudy (anonymous):

@jollysailorbold

OpenStudy (anonymous):

I SAID D ?

OpenStudy (anonymous):

that was not what was asked of you.... please answer her question...

OpenStudy (anonymous):

2 & 6 ?

OpenStudy (diyadiya):

Well, By the way how did you get the answer?

OpenStudy (diyadiya):

2& 6? hmm i said "two numbers whose product is -32 and sum is 14 " 2 x 6= 12 not -32 2 + 6 = 8 not 14

OpenStudy (anonymous):

7*2

OpenStudy (diyadiya):

what is 7 * 2 = ?

OpenStudy (anonymous):

14

OpenStudy (diyadiya):

But product should be -32

OpenStudy (anonymous):

-16*2

OpenStudy (diyadiya):

Yes! now add them and see if their sum is 14 or not

OpenStudy (anonymous):

YES

OpenStudy (diyadiya):

-16 + 2 = ?

OpenStudy (anonymous):

14

OpenStudy (diyadiya):

Nope

OpenStudy (anonymous):

-14

OpenStudy (diyadiya):

-16 + 2 =-14

OpenStudy (diyadiya):

Yes but we need 14 ,so try another

OpenStudy (anonymous):

I DONT UNDERSTAND WHAT YOUR TRYING TO SAY !

OpenStudy (diyadiya):

Ok letme explain

OpenStudy (diyadiya):

x^2 + 14x – 32 this is of the form ax^2 + bx + c (here a=1, b= 14, c=-32 ) while factoring we have to find tw numbers who'se product is 'a*c' and sum is 'b' here a =1 & c=-32 so product should be -32 b=14 so sum should be 14

OpenStudy (diyadiya):

and you gave me the numbers -16 & 2 their product is -32 (that's right) But their sum is not 14

OpenStudy (diyadiya):

So, lets take the numbers 16 & -2 what is the product & sum here?

OpenStudy (anonymous):

tanga letter c is the answer!

OpenStudy (anonymous):

@Ceejay @Diyadiya was helping answer this. We do not encourage just giving answers on this site, as is outlined in CoC. We are teaching people, not doing their homework.